# ManiSkill3: GPU Parallelized Robotics Simulation and Rendering for Generalizable Embodied AI

### v1

@@ Appendix B Simulation and Rendering Benchmarking
We carefully analyze the performance of ManiSkill3’s parallel simulation and simulation+rendering performance compared to Isaac Lab. Currently we only have accurate results for a simple cartpole environment. We compare against the latest Isaac Lab v1.2.0 released on Sept. 20, 2024.
B.1 Setup
@@
Figure 13: Simulation FPS against number of parallel environments without rendering. Annotated numbers on top of data points indicate the GPU memory usage.
B.2 Simulation Only Benchmark Results
Testing on the cart pole task, we see that ManiSkill3 always runs a little faster and uses about 1.5x to 2x less GPU memory in Figure 13. We note that while ManiSkill3 appears faster here, one could make a many simulation specific optimizations (simplified collisions meshes, tuned physics solver configurations etc.) to increase simulation speed and trade off simulation accuracy. We specifically create documentation/tutorials on the various tricks/optimizations a user can perform to improve simulation speed and/or fidelity from the environment object/robot models to simulation configurations.
B.3 Simulation+Rendering Benchmark Results
We ablate on a number of aspects of visual data collection where we simulate and render an environment. We record the FPS of taking 1000 random actions in the environment and fetching the 1000 visual observations. Generally in the ablations we scale up one of environments, camera sizes, or number of cameras until the benchmark runs out of GPU memory.
Overall Isaac Lab shows about 2-3x higher GPU memory usage compared to ManiSkill3, with the minimum amount of GPU memory taken up by just environments with cameras enabled being 4.8GB compared to 1.7GB in ManiSkill3.
B.3.1 Ablation on Realistic Camera Settings
We test on two realistic camera resolutions/setups based on the setup of the Open-X dataset (Collaboration et al., 2023) and the Droid dataset (Khazatsky et al., 2024). Many datasets in Open-X like the Google RT datasets have a single 640x480 resolution RGB camera, and the Droid dataset uses three 320x180 resolution RGB cameras. The results are shown in Figures 14 and 15. In both cases ManiSkill3 has about 2-3x better GPU memory efficiency and runs about 2x faster. Note that Isaac Lab’s rendering output is a bit different from ManiSkill3, we share some qualitative examples in Appendix B.4. We further note that both ManiSkill3 and Isaac Lab support more photorealistic, high quality renders with full ray-tracing that is not parallelized. Example ManiSkill3 high-quality ray-traced renders can be seen in Figure 1.
@@ B.3.1 Ablation on Realistic Camera Settings
B.3.2 Ablation on Camera Size
We ablate on square camera sizes from a large 512x512 resolution to a small 128x128 resolution shown in Figures 16 17 18. We observe that Isaac Lab runs at most about 1.2x faster when there are a high number of parallel environments with small camera resolutions compared to ManiSkill3. At larger camera resolutions ManiSkill3 outperforms up to 2x in speed and 3x in GPU memory usage. Notably, ManiSkill3 always outperforms about 2-4x in speed and 2-3x in GPU memory usage for smaller number of parallel environments. While Isaac Lab is fast at smaller resolutions, we note that some manipulations tasks are fairly impractical and impossible to solve at small resolutions. Moreover memory efficiency is absolutely critical for RL applications that can remain fast by keeping large replay buffers on the GPU instead of using CPU memory. For this reason all visual RL baselines in ManiSkill3 typically do not use more than 256 parallel environments as RL rollouts only run marginally faster with more environments beyond that point (for both Isaac Lab and ManiSkill3) but GPU memory use worsens a lot more.
@@ B.3.2 Ablation on Camera Size
B.4 Qualitative Rendering Results
To the best of our capabilities we try to make the benchmarked environments look as similar as possible in terms of visuals. It is not possible to keep everything the same given fundamental differences in the parallel rendering system of Isaac Lab compared to ManiSkill3 or easily quantify the rendering differences. Section 3.2 details a little bit about the differences. Figure 4 in the main paper shows the cartpole RGB+Depth rendering results. Figures 19 and 20 shows RGB renders of a environment with a Franka Panda arm with different resolutions.

### v2

@@ XI Simulation and Rendering Benchmarking
We carefully analyze the performance of ManiSkill3’s parallel simulation and simulation+rendering performance compared to Isaac Lab. Currently we only have accurate results for a simple cartpole environment. While one can try and compare more complex environments it is difficult to align things perfectly. We compare against the Isaac Lab v1.2.0 which was first version with parallel rendering support that was released a few months after we initially released our framework.
XI-A Setup
@@
Figure 30: Simulation FPS against number of parallel environments without rendering. Annotated numbers on top of data points indicate the GPU memory usage. ManiSkill speed for this task is on par with Isaac Lab.
XI-B Simulation Only Benchmark Results
Testing on the cart pole task, we see that ManiSkill3 runs a little faster and uses about 1.5x to 2x less GPU memory in Figure 30. We note that while ManiSkill3 appears faster here, one could make many simulation specific optimizations (simplified collisions meshes, tuned physics solver configurations etc.) to increase simulation speed and trade off simulation accuracy. We further note it is difficult to make truly apples-to-apples comparisons between simulators and that this is simply just one data point and may not extend to other environments. We specifically create documentation/tutorials on the various tricks/optimizations a user can perform to improve simulation speed and/or fidelity from the environment object/robot models to simulation configurations.
XI-C Simulation+Rendering Benchmark Results
We ablate on a number of aspects of visual data collection where we simulate and render an environment. We record the FPS of taking 1000 random actions in the environment and fetching the 1000 visual observations. Overall Isaac Lab shows about 2-4x higher GPU memory usage compared to ManiSkill3, with the minimum amount of GPU memory taken up by just environments with cameras enabled being 4.8GB compared to 1.7GB in ManiSkill3.
XI-C1 Ablation on Realistic Camera Settings
We test on two realistic camera resolutions/setups based on the setup of the Open-X dataset [14] and the Droid dataset [28]. Many datasets in Open-X like the Google RT datasets have a single 640x480 resolution RGB camera, and the Droid dataset uses three 320x180 resolution RGB cameras. The results are shown in Figures 31 and 32. In both cases ManiSkill3 has about 2-4x better GPU memory efficiency and runs about 2x faster. Note that Isaac Lab’s rendering output is a bit different from ManiSkill3, we share some qualitative examples in Appendix XI-D. We further note that both ManiSkill3 and Isaac Lab support more photorealistic, high quality renders with full ray-tracing that is not parallelized. Example ManiSkill3 high-quality ray-traced renders can be seen in Figure 1.
@@ XI-C1 Ablation on Realistic Camera Settings
XI-C2 Ablation on Camera Size
We ablate on square camera sizes from a large 512x512 resolution to a small 128x128 resolution shown in Figures 33 34 35. We observe that Isaac Lab runs at most about 1.25x faster when there are a high number of parallel environments with small camera resolutions compared to ManiSkill3. At larger camera resolutions ManiSkill3 outperforms up to 2x in speed and 4x in GPU memory usage. Notably, ManiSkill3 always outperforms about 2-4x in speed and 2-3x in GPU memory usage for smaller number of parallel environments. While Isaac Lab is fast at smaller resolutions, we note that some manipulations tasks are fairly impractical and impossible to solve at small resolutions. Moreover memory efficiency is absolutely critical for RL applications that can remain fast by keeping large replay buffers on the GPU instead of using CPU memory. For this reason all visual RL baselines in ManiSkill3 typically do not use more than 256 to 1024 parallel environments as RL rollouts only run marginally faster with more environments beyond that point (for both Isaac Lab and ManiSkill3) but GPU memory use worsens a lot more.
@@ XI-C2 Ablation on Camera Size
XI-D Qualitative Rendering Results
To the best of our capabilities we try to make the benchmarked environments look as similar as possible in terms of visuals. It is not possible to keep everything the same given fundamental differences in the parallel rendering system of Isaac Lab compared to ManiSkill3 or easily quantify the rendering differences. Section III-B details a little bit about the differences. Figure 3 in the main paper shows the cartpole RGB+Depth rendering results. Figures 36 and 37 shows RGB renders of a environment with a Franka Panda arm with different resolutions.
@@ XI-D Qualitative Rendering Results
XII VR Teleoperation
ManiSkill3 provides VR support for all mainstream VR devices by implementing the OpenVR client protocol designed by Steam. Specifically, ManiSkill3 receives camera intrinsic and extrinsic parameters of the head-mounted display, VR controllers’ poses, and the operator’s hand and wrist poses if supported by the hardware; and it sends stereo video streams at 4K resolution for OpenVR to display in the headset. Under the hood, OpenVR communicates with SteamVR and ALVR which translate hardware-dependent VR implementation into the unified OpenVR client protocol.
The VR feedback loop runs at 60 Hz which is crucial for smooth user experience, while an asynchronous action translator translates sensed VR poses to robot actions at 20 Hz, so we can tolerate computationally intensive translation algorithms.
Figure 38: Overview of our VR teleoperation system. The operator controls the robot’s arm and hand through real-time tracking of wrist and hand poses, while receiving stereo video feedback. The VR devices stream human pose data to a server, which retargets and sends joint commands to the robot.
Overview for our VR teleoperation system is shown in Figure 38. In a simulation setup, the stereo video is rendered based on simulated environments. For real-world robot teleoperation, the stereo video is generated from point clouds captured by depth cameras. We utilize the SAPIEN engine to achieve high-speed rendering of these high-resolution stereo videos. To simplify the complex SteamVR setup, we provide a Docker image for a smoother and faster setup process for users.
Figure 39: Demonstration of VR teleoperation across various tasks and robotic setups. Top row: Sequential images of a Panda arm performing precise insertions – charging connector insertion (left) and peg insertion (right) – showcasing accurate teleoperation for tasks requiring high precision. Middle row: Sequence of dual XArm7 robots with grippers demonstrating the grasping and unscrewing of a bottle, highlighting multi-arm coordinated teleoperation. Bottom row: Various teleoperation setups and manipulations, including (from left to right) XArm7 with a gripper, a floating Delto 3-finger hand, XArm7 with Inspire hand, dual XArm7 with Ability hand, a floating Allegro hand, and real-time teleoperation of an XArm7 with a gripper grasping objects.
XII-A Control
The human-motion-to-robot-action conversion system consists of three main modules:
1) Arm Control Module, which converts human wrist poses into robot arm joint positions. However, directly mapping the absolute orientation and position of wrist poses to robot’s end-effector poses can lead to strange behaviors. For example there can be a mismatch between the coordinate frames of human wrist poses and the robot’s end-effector frame as illustrated in Figure 40. To address this, we provide configurations out-of-the-box for the most common robotics arms. Additionally, we offer a GUI and tools to assist users in quickly computing the transformation matrix for their customized robots.
Our system utilizes a modified version of the Closed-loop Inverse Kinematics (CLIK) algorithm, implemented with the Pinocchio library [6, 7], to calculate the joint angles of the robot’s arm.
The approach uniquely addresses the problem of inverse kinematics (IK) for two coupled end-effectors simultaneously, which is critical in dual-arm robots with a sliding joint at the shoulders.
For a robot with $n$ joints and two coupled end-effectors $EE_{i}$, we aim to find joint position $q$ that minimizes the pose error for both end-effectors relative to their respective target poses $T_{i}$. Solving IK independently for two end-effectors suffers from large errors. Alternatively, iterative approaches usually result in slow convergence. Our solution extends the standard CLIK by computing a concatenated Jacobian that accounts for both end-effectors simultaneously. This enables faster and more accurate optimization of dual-arm configurations.
Additionally, the system allows users to control the movement of specific joints during coordinated motion. Instead of using a binary mask (0 for no movement, 1 for full movement) to constrain joints, we employ a soft mask with values between 0 and 1. A lower mask value reduces motion in the corresponding joint, leading to a solution that minimizes undesired movement. To ensure smooth arm motions, we apply an SE(3) group filter to the input end-effector poses before the IK computations.
2) Hand Control Module, which translates human finger poses into corresponding robot hand joint positions. Following [41, 11, 15], we formulate the hand motion retargeting process as an optimization problem. The objective function for this optimization is defined as follows:
$\min_{q_{t}}\sum_{i=0}^{N}\|\alpha^{i}v_{t}^{i}-f_{i}(q_{t})\|^{2}+\beta\|q_{t}-q_{t-1}\|,$ (1)
where $q_{t}$ denotes the robot hand joint positions at time $t$, $v_{t}^{i}$ is the $i$-th keypoint vector of human hand, and $f_{i}(q_{t})$ gives the corresponding $i$-th keypoint vector of robot hand using forward kinematics with joint positions $q_{t}$. The scaling factor $\alpha_{i}$ compensates for the differences in hand size between the human and robot hands and treat each $i$-th keypoint differently as thumb finger size and pinky finger size can vary a lot, $\beta$ weights the regularization term to ensure temporal consistency between consecutive joint positions. The optimization is implemented by NLopt solver [41].
For dexterous robot hands, we map vectors from human hand fingertips to palm base to corresponding vectors on the robot hand. In certain cases, such as with the Inspire hand, additional vector mappings are used to improve motion accuracy–for example, an extra vector from the thumb metacarpophalangeal joint to the thumb tip is employed. For grippers, we use a single vector optimization between the human thumb and index fingertips, which corresponds to the gripper’s upper and lower ends. This allows intuitive control over the gripper’s opening and closing motions through simple pinching gestures with the operator’s index finger and thumb [11].
We fine-tune configurations for several common robots and provide calibration tools to assist users in adapting the system to their custom robots.
3) Controller Control Module, which is also used to control gripper motion, allowing simple and effective control of grippers. Though Apple Vision Pro does not include a physical controller, our system still supports controller-based input. By clipping the controller, users can trigger the closing action of the gripper, enabling intuitive and responsive control during operation.
Figure 40: Illustration of frame mismatch between human wrist poses and the robot’s end-effector frame. Left: human wrist frame, Right: Inspire hand frame as an example.
XII-B Sim-to-Real Interface
Our system employs a unified interface for both simulation and real-world setups by aligning the robot’s end-effector with the absolute position of human hand. This alignment helps teleoperators better understand spatial information. Previous approaches [41, 11, 15] struggled to achieve spatial alignment between the human hand and the robot hand, forcing operators to compensate for parallax effects, which made the teleoperation experience less intuitive.
A key challenge arises when attempting to align the human hand with the robot hand inside a VR headset, especially since the real robot may be spatially displaced in the physical world. To solve this, we project the point clouds captured by depth cameras positioned around the real robot into the VR headset. The camera poses are calibrated using EasyHec [8].
This setup ensures that both the simulation environment and the real-world point cloud are aligned in a ”digital twin” manner as illustrated in Figure 41. Though we do not require the visual textures to match the real world, critical elements such as the robot’s position, forward and inverse kinematics, and control interface must be aligned. This alignment allows the same human control signals to produce identical robot actions in both the simulator and the real-world environment. Consequently, teleoperation in the real world becomes as intuitive and consistent as it is in simulation.
Figure 41: Illustration of spatial alignment of simulation and real-world environments. Left: simulation environment, where robot control are tested, Middle: real-world teleoperation setup, Right: point cloud captured from the depth cameras accurately aligned with the virtual robot in the simulation, demonstrating a ”digital twin” setup.